On July 1, 2017, Lilly paid $90,000 for 450 shares of Jennifer Corp common stock. Lilly...

  1. On July 1, 2017, Lilly paid $90,000 for 450 shares of Jennifer Corp common stock. Lilly received a nontaxable stock dividend of 50 new common shares on December 15, 2019. On December 20, 2019, Lilly sold the 50 new shares for $11,000. How much should Lilly report in her 2019 return as Long Term Capital Gain?
  1. $0
  2. $1,000
  3. $2,000
  4. $11,000
Answer #1

Solution:

Calculation of long term capital gain = sales - cost

sales of 50 share is 11,000

Cost of each share = 90,000 / 450

=200

cost of 50 shares = 50*200

=10,000

Calculation of long term capital gain = sales - cost

= 11,000 - 10,000

=1,000

Option B)$1,000 is correct